A gas is filled in a container of volume v at 121 degrees celsius to what temperature should it be heated in order that 1/4th off the gas may escape out of the vessel

Grade:12

2 Answers

Arun
6 years ago
Ideal gas equation is PV=nRT. Initially T=121+273=394 k so PV=nR(394)....(1) After since 1/4 of the gas is escaped so n`=n-1/4n =3n/4 so PV=(3n/4)RT`....(2)now equating equations we get nR394=3n/4RT` we get T`=525k=252`c here v remains same because the remaining gas also occupies whole of the container volume.(since gas have infinite expandability).
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Ideal gas equation is PV=nRT.
Initially T = 121 + 273 = 394k
so PV = nR(394)....(1)
After since 1/4 of the gas is escaped so
n`=n-1/4n =3n/4
so PV=(3n/4)RT`....(2)
now equating equations we get
nR394=3n/4RT`
we get T`=525k
=252`c
here v remains same because the remaining gas also occupies whole of the container volume.(since gas have infinite expandability).

Thanks and Regards